# Hybrid Attention Network: An efficient approach for anatomy-free landmark detection

### v1

@@ 4.4.1 Head dataset results
Figure 4: Visualizations of various methods on the ISBI2015 dataset. The red points represent the
predicted landmarks, while the green points correspond to the ground truth labels. Local details are
provided below for a clearer comparison of the results. The MRE value is shown in the top-left corner
for reference.
@@
Following the approach in the CeLDA [41], we further report results separately for adolescent and adult data. Our method demonstrates exceptional performance on the adolescent subset, achieving an MRE of 0.85px, which is 15% reduction compared to the SOTA result. Additionally, SDR improves across all thresholds, with the SDR within 2mm increasing from 89.93% to 91.33%. When compared to other general methods such as U-Net, GU2Net, and SRLD-Net, our method consistently shows superior SDR at all detection thresholds, with more pronounced improvements. As illustrated in Fig. 6, there are clear anatomical differences between the left-side adolescent image and the right-side adult image, yet our method accurately predicts landmarks in both cases.
@@ 5 Conclusion
This paper presents the Hybrid Attention Network(HAN) for accurate and efficient anatomical landmark detection in medical images. This novel architecture combines CNNs and Transformers, using a dynamic sparse attention mechanism (BiFormer with Bi-Level Routing Attention) to efficiently handle high-resolution images. The BiFormer captures global context, while Convolutional Attention Blocks (CABs), enhanced by CBAM, refine local features. A Feature Fusion Correction Module (FFCM) integrates multi-scale features, preventing resolution loss. Extensive experiments across five diverse datasets demonstrate state-of-the-art performance, exceeding existing methods in accuracy, robustness, and efficiency. The HAN offers a promising framework for various medical image analysis tasks. Future work will explore 3D image applications and further optimize the sparse attention mechanism. This method has the potential to significantly improve clinical practice through more accurate and efficient image-guided procedures.

### v2

@@ 4.4.1 Head dataset results
Figure 4: Visualizations of various methods on the ISBI2015 dataset. The red points represent the predicted landmarks, while the green points correspond to the ground truth labels. Local details are provided below for a clearer comparison of the results. The MRE value is shown in the top-left corner for reference.
@@
Following the approach in the CeLDA [46], we further report results separately for adolescent and adult data. Our method demonstrates exceptional performance on the adolescent subset, achieving an MRE of 0.85px, which is 15% reduction compared to the SOTA result. Additionally, SDR improves across all thresholds, with the SDR within 2mm increasing from 89.93% to 91.33%. When compared to other general methods such as U-Net, GU2Net, and SR-UNet, our method consistently shows superior SDR at all detection thresholds, with more pronounced improvements. As illustrated in Fig. 6, there are clear anatomical differences between the left-side adolescent image and the right-side adult image, yet our method accurately predicts landmarks in both cases.
@@ 5 Conclusion
This paper presents the Hybrid Attention Network(HYATT-Net) for accurate and efficient ALD in medical images. This novel architecture combines CNNs and Transformers, using a dynamic sparse attention mechanism (BiFormer with Bi-Level Routing Attention) to efficiently handle high-resolution images. The BiFormer captures global context, while Attention Residual Blocks (ARMs), enhanced by CBAM, refine local features. A Feature Fusion Correction Module (FFCM) integrates multi-scale features, preventing resolution loss. Extensive experiments across five diverse datasets demonstrate state-of-the-art performance, exceeding existing methods in accuracy, robustness, and efficiency. The HYATT-Net offers a promising framework for various medical image analysis tasks. Future work will explore 3D image applications and further optimize the sparse attention mechanism. This method has the potential to significantly improve clinical practice through more accurate and efficient image-guided procedures.